OpenStudy (anonymous):

solve equation -5(b+7)=4(3b+4) how do you do it?

jimthompson5910 (jim_thompson5910):

what do you get when you distribute -5(b+7)

OpenStudy (anonymous):

so just distributing? -5b+35

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but keep in mind that the negative distributes as well

OpenStudy (anonymous):

oh ok so -5b-35

jimthompson5910 (jim_thompson5910):

good, what about 4(3b+4)

OpenStudy (anonymous):

12b+16

jimthompson5910 (jim_thompson5910):

so we have this so far -5(b+7)=4(3b+4) -5b-35=12b+16

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

combine like terms

jimthompson5910 (jim_thompson5910):

are there any like terms?

jimthompson5910 (jim_thompson5910):

think of each side as separate things for now (you can't combine across the equal sign)

OpenStudy (anonymous):

-5b -35 = 12b+16 -5b -12b = 16+35 -17b = 51 b = -3

jimthompson5910 (jim_thompson5910):

very good, you nailed it

OpenStudy (anonymous):

that's what i was saying get the like terms on each side

OpenStudy (anonymous):

then combine

jimthompson5910 (jim_thompson5910):

ah i gotcha

OpenStudy (anonymous):

yw

jimthompson5910 (jim_thompson5910):

I thought you meant combine the like terms in -5b-35=12b+16

jimthompson5910 (jim_thompson5910):

but there aren't any because they are separated

jimthompson5910 (jim_thompson5910):

but you got it though

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

oh i was reading the wrong post...thought you were the OP

jimthompson5910 (jim_thompson5910):

sorry about that

OpenStudy (anonymous):

like 2 ( 4n+7)= -2(n+8) +5n is equal to -6

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

yes, n = -6 is the solution

OpenStudy (anonymous):

ok i got it thanks

jimthompson5910 (jim_thompson5910):

np